OpenStudy (anonymous):

9. Write the standard equation for the circle center (–6, –8), that passes through (0, 0).

OpenStudy (anonymous):

(x-h)^2+(y-k)=r^2 whare h,k is center of the cirlce.

OpenStudy (anonymous):

to find radious Just use distance formula using point -6,-8 and 0,0

OpenStudy (anonymous):

so it be (x-(-6))^2+(y-(-8))^2= sqrt((-6-0)^2+(-8+0)^2)

OpenStudy (anonymous):

well it be (x-(-6))^2+(y-(-8)^2= 10^2 right?

OpenStudy (anonymous):

there is a mistake in rhs side it is (sqrt((-6-0)^2+(-8+0)^2))^2